The 2017 balance sheet of Kerber's Tennis Shop, Inc., showed long-term debt of $5.4 million, and the 2018 balance sheet showed long-term debt of $5.6 million. The 2018 income statement showed an interest expense of $175,000. During 2018, the company had a cash flow to creditors of –$25,000 and the cash flow to stockholders for the year was $80,000. Suppose you also know that the firm’s net capital spending for 2018 was $1,390,000, and that the firm reduced its net working capital investment by $73,000. What was the firm’s 2018 operating cash flow, or OCF? (Enter your answer in dollars, not millions of dollars, e.g., 1,234,567.)
Answer #1

1) Cash flow from assets

Cash flow from assets = Cash flow to Creditors+ Cash flow to Stockholders

= -25,000+80,000

= $55,000

2) Operating Cash Flow (OCF)

Cash flow from assets = OCF - Change in Net Working Capital - Net Capital Spending

55,000 = OCF + 73,000 - 1,390,000

OCF = 55,000-73,000+1,390,000

= $1,372,000